OpenStudy (anonymous):

Molly finds two sticks and wants to build a kite for her little brother. The sticks are 12 inches and 24 inches long, and she plans on using them as the diagonals of the kite. Molly will need a minimum of _____ square inches of fabric to make the kite.

jimthompson5910 (jim_thompson5910):

Hint: If p and q are the diagonals, then the area of the kite is A = (p*q)/2

OpenStudy (anonymous):

12*24=288/2=144?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

yay!!!!!! i did it by head too:D

jimthompson5910 (jim_thompson5910):

lol very nice

OpenStudy (anonymous):

thank you:)

jimthompson5910 (jim_thompson5910):

you're welcome